OpenStudy (anonymous):

half life of carbon-14 is 5730 years. all radioactive decay is first order. measurments taken indicate that scrolls contain 79.5% of carbon-14. how old are these scrolls? I derived the rate law to get half life so I could solve for k. but now I don't know what to do

OpenStudy (unklerhaukus):

\[N(t)=N_0e^{-\lambda t}\\ \frac{N(t)}{N_0}=e^{-\lambda t}\\ t =-\frac1\lambda\ln\frac{N(t)}{N_0}\] \[\lambda = 1/\tau\\\qquad\qquad\qquad\qquad\tau=\frac{t_{1/2}}{\ln 2}\\ \lambda=\frac{\ln 2}{t_{1/2}}\] \[t=\frac{-t_{1/2}}{\ln 2}\ln\frac{N(t)}{N_0}\]

OpenStudy (perl):

You can use the equation y = A * (1/2) ^ (t/5730) where A is the initial amount of carbon-14 in grams and y is the amount remaining after t years We are given that there 79.5% of the initial amount remaining. .795*A = A * ( 1/2)^(t / 5730) you can divide both sides by A .795 = ( 1/2) ^( t / 5730) solve for t ln( .795) = ln( 1/2) ^( t / 5730) ln(.795) = t/5730 * ln(1/2) ln(.795)/ln(1/2) = t/5730 5730 * ln(.795)/ln(1/2) = t

OpenStudy (perl):

Can you solve it now?